If a trust fund amount to#57,000 was invested five years ago at 15 percent compounded annually.find the amount in the fund now
1
Expert's answer
2021-11-12T03:55:03-0500

"A=P(1 +\frac {r} {100}) ^{n}"


"A=57,000(1 +0.15)^5"


"A=57,000(1.15) ^5"


"A=57,000\u00d72.01"


"A=\$114,570"